Hey guys,
I really need some help with this question please.
Two towns are 1100 miles apart. A group of hikers start from each town and walk the trail toward each other. They meet after a total hiking time off 200 hours. If one group travels 1/2 mile per hour faster than the other group, find the rate of each group for the slower and faster group.
Thank you

Answer by rfer(16322)

1100/200=5.5 miles per hr
x+x+0.5=5.5
2x=5.5-0.5
2x=5
x=5/2
x=2.5 mph
x+0.5=3 mph